Write each logarithmic expression as a single logarithm.

24. log 2 + 3 log 1 2
25. log a - log ab

Answer by Alan3354(69443)

24. log 2 + 3 log 1 2
= log(2) + log(12^3)
= log(2*12^3)
= log(3456)
-------------
25. log a - log ab
= log(a/ab)
= log(1/b)
= -log(b)